OpenStudy (anonymous):

Which value for x makes the sentence true? 3x = -2x + 25

OpenStudy (anonymous):

3X = -2X + 25 +2X +2X 5X = 25 /5 /5 X = 5

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

No problem :)